OpenStudy (anonymous):

If an object is thrown directly upwards (initial speed 3m/s) what will its maximum height be?

OpenStudy (anonymous):

oh, and the only other force acting on it is gravity (9.8m/s)

ganeshie8 (ganeshie8):

there are several ways to do this - my fav is conservation of energy : inital PE + KE = final PE + KE

OpenStudy (anonymous):

what is PE and KE

ganeshie8 (ganeshie8):

good question, PE = Potential Energy KE = Kinetic Energy

ganeshie8 (ganeshie8):

since initially there was no PE, and at maximum height there wont be any KE, the equation above simplifies to : initial KE = final PE

ganeshie8 (ganeshie8):

\(\large \dfrac{1}{2}m v_i^2 = mgh\)

ganeshie8 (ganeshie8):

\(\large \dfrac{1}{2} v_i^2 = gh\) \(\large \dfrac{v_i^2}{2g} = h\)

ganeshie8 (ganeshie8):

plugin the values and evaluate ^

OpenStudy (anonymous):

so \[v _{i} \] is 2m/s

ganeshie8 (ganeshie8):

you're given \(\large v_i = 3m/s\) right ?

OpenStudy (anonymous):

oh, sorry...yep I meant that

OpenStudy (anonymous):

haha!! If I plug those values in I get h=0.2093 which is very small?

OpenStudy (anonymous):

wait... h=0.469

OpenStudy (anonymous):

that's still really small...

ganeshie8 (ganeshie8):

throw the stone bit faster if u want it go high :)

OpenStudy (anonymous):

oh...so that's right?

OpenStudy (anonymous):

can I also ask in the first equation you gave what does m mean?

ganeshie8 (ganeshie8):

yep u get that much only for that speed lol :P

OpenStudy (anonymous):

\[mass \times gravity \times height\]

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

No problem